OpenStudy (anonymous):

Let a and b be integer with \[ a\ge 0\\ b \ge 0\\ 2^n a + b \text { is a perfect square for any natural number } \, n \, \] Show that a=0.

OpenStudy (anonymous):

hello doc... :)

OpenStudy (anonymous):

Hi @dpalnc

OpenStudy (anonymous):

hello @eliassaab

OpenStudy (shubhamsrg):

is it correct? because for ,,let n=5 and a=1,,b=32 gives 64,,which is a perfect square..

OpenStudy (anonymous):

It is a perfect square for every n and not only for a particular n.

OpenStudy (shubhamsrg):

what i wanted to say is that when a not equals 0,,then also its a perfect square..

OpenStudy (shubhamsrg):

and the ques asks to prove a=0

myininaya (myininaya):

So it is \[2^na+b ?\] or \[2^{na+b}\]? I just wanted to be sure you just meant to raise just the n

OpenStudy (anonymous):

\[ 2^n \, a\, + b \]

OpenStudy (shubhamsrg):

but as i said above,,when a not equals 0,,then also the eqn can be a perfect sq..

OpenStudy (anonymous):

You have to show that if a is not zero and \[ 2^n \, a + b \] is a perfect square for every n, then you must conclude a contradiction. ( for every n) and not only for a particular n.

OpenStudy (shubhamsrg):

ohh..not completely my fault though,,in the ques,,you wrote show a=0!! ;)

OpenStudy (anonymous):

But I said " is a perfect square for any natural number n Show that a=0. "

OpenStudy (shubhamsrg):

not getting it!! o.O bah!1 let it be..seems like not my cup of tea..

OpenStudy (anonymous):

Clearly the method to use is proof by contradiction, find a contradiction for the assumption that \(a>0\). From there, though, I am not yet sure where to go.

OpenStudy (dumbcow):

lets see if you look at cases where n is even then 2^n is itself a perfect square when n is odd, then 2^n is not a perfect square but a,b must satisfy all n n is even, let a=1, then b must be 0 n is odd, and a=1, b=0 --> result is not a perfect square n is even, let a = 0, then b must be some constant perfect square n is odd , a=0, b = constant --> result perfect square

OpenStudy (experimentx):

Let n=1,a=1, b=2 we have 2^1*1 + 2 = 4, perfect square but a=1 != 0 Not sure if i am getting the question right!!!

OpenStudy (anonymous):

@dumbcow is correct, but that is only a contradiction for a=1, not for a>0. We need a contradiction for all a>0.

OpenStudy (dumbcow):

yeah i need to generalize it but thats the right idea i believe

OpenStudy (anonymous):

Yes, definitely right place to start. I think we can do a case-by-case with n=odd or n=even and a>0 to find a contradiction in general.

OpenStudy (anonymous):

I dont understand the question :/ What is the question asking?

OpenStudy (anonymous):

\(2^na + b\) is perfect square only when \(a=0\)?

OpenStudy (dumbcow):

correct for all n

OpenStudy (anonymous):

can we do it by induction? assuming the \(2^na + b\) to be prefect for \(a \neq 0\).

OpenStudy (experimentx):

seems like i misunderstood question!!

OpenStudy (anonymous):

perfect square* For \(n=1\), \(a+b\) is perfect square. For \(n=k\), Let \(2^ka + b\) be perfect square. Hmm I dont think it's right :/

OpenStudy (experimentx):

b has to be a perfect square ..

OpenStudy (anonymous):

Let \(2^ka+b=q^2\). \[2^{k+1}a + b = 2^ka + 2^ka + b = q^2 + 2^ka\]

OpenStudy (anonymous):

Hmm what's the condition for Pythagorean triplet? but before that, I will need to show if \(2^ka\) is perfect square or not for \(a\neq 0\).

OpenStudy (dumbcow):

are you trying to show a contradiction when a != 0 by using induction ?

OpenStudy (anonymous):

yeah, but i am probably not getting anywhere.

OpenStudy (anonymous):

If \(k\) is even, a has to be perfect square and if \(k\) is odd then a must be odd powers of 2 for \(2^ka\) to be perfect square.

OpenStudy (anonymous):

hmm for \(2^ka + b = q^2\) to be valid, \(2^ka\) and \(b\) have to be perfect square, other than being pythagorean triplets. For \(k+1\), \(2^{k+1}a + b\) can't be perfect square if \(a\neq 0\). Because: "If \(k\) is even, \(a\) has to be perfect square and if \(k\) is odd then \(a\) must be odd powers of 2 for \(2^ka\) to be perfect square."

OpenStudy (anonymous):

Is this good enough?

OpenStudy (anonymous):

No. Here a hint: First show that b cannot be zero. Then show that the hypothesis and a not zero implies the existence of two integers u and v such that \[ u^2 < v^2 < (u+1)^2 \] Which is impossible.

OpenStudy (kinggeorge):

Here's the way I'm thinking about this. Suppose \(b=0\), then either \(2a\) or \(4a\) is not a perfect square (for \(a\neq0\)). So let \(b\neq0\). Also, at some point, we know that \[2^na\leq b< 2^{n+1}a\]If \(b=2^na\) then \(2^na+b=2^{n+1}a\) and we're back to square one. If \(b\neq 2^{n}a\), then we have that\[2\cdot2^{n}a<2^na+b<3\cdot2^na\]If we write \(u=2^na\) we have the relation \[2 u>u+b <3u\]At some points, this must be false since \(u\) varies. So that leaves us with two possibilities. 1. \(b=0\). If true, \(a=0\) by the first part. 2. \(b\neq0\). If this is true, we want \(u=2^na\) to not vary. Hence, \(a=0\). QED

OpenStudy (anonymous):

Should 2u>u+b<3u be 2u<u+b<3u

OpenStudy (anonymous):

Let \[ u_n^2 = 2^{n+2} + b, \quad b \ne 0,\, a\ne0\\ \] Choose n so that \[ u_n > \frac 3 2 b \\ (u_n+1)^2 = u_n^2 + 2u_n +1 > u_n^2 +3b =\\2^{n+2}a + 4 b = 4( 2^n a + b) = v^2 > 2^{n+2}a + b= u_n^2 \] if we put \[ u = u_n\\ \text { then }\\ (u+1)^2 > v^2 > u^2 \] Contradiction.

OpenStudy (kinggeorge):

Yes, \(2u>u+b<3u\) should be \(2u<u+b<3u\). Thanks for catching that.